OpenStudy (scorcher219396):

Evaluate: lim x-> 0 sinx/(x+tanx)

OpenStudy (scorcher219396):

@SithsAndGiggles

OpenStudy (anonymous):

Quick question, have you learned L'Hopital's rule? It's not required, but it's one way to go about this.

OpenStudy (scorcher219396):

Nope, not yet. (only ch 3 right now, just starting to take derivatives)

OpenStudy (anonymous):

Another way would be to make use of the fact that near values of \(x=0\), you have \(\sin x\approx\tan x\approx x\). This is the simplest way.

OpenStudy (anonymous):

Yet another way is to use identities and known limits like before. Your choice.

OpenStudy (scorcher219396):

Can you show me the way with identities and limits since I'm assuming that's how the textbook wants?

OpenStudy (anonymous):

Sure. \[\large\begin{align*} \lim_{x\to0}\frac{\sin x}{x+\tan x}&=\lim_{x\to0}\frac{\sin x}{x+\frac{\sin x}{\cos x}}\\\\ &=\lim_{x\to0}\frac{1}{\frac{x}{\sin x}+\frac{1}{\cos x}}&\text{divide num and den by }\sin x\\\\ &=\dfrac{\displaystyle\lim_{x\to0}1}{\displaystyle\lim_{x\to0}\frac{x}{\sin x}+\lim_{x\to0}\frac{1}{\cos x}} \end{align*}\]

OpenStudy (anonymous):

*divide num and den by \(\sin x\)*

OpenStudy (scorcher219396):

Ok got it! Thanks!